OpenStudy (laddiusmaximus):

a ramp 12ft long rises to a platform that is 10ft off the ground. find x, the elevation of the ramp. round answer to the nearest hundredth of a degree

OpenStudy (anonymous):

tan(x) = opposite/ adjacent tanx = 10/12 x = arctan(5/6) x =39.81 degrees